I need help with these 3 problems. Thank you!
A shop sells compact disks for $15 a piece. If each disk costs $6 to the shop owner how much profit percent does the shop owner earn?
A computer company gives a discount of 10% on a new piece of software. If it makes a profit of 4%, what percent does the marked price exceed the cost price?
What single discount is equivalent to two successive discounts of 15% and 7%?

Answer by edjones(8007)

Let p=profit
6p=15
p=15/6
=5/2=2.5=250%
.
Ed